OpenStudy (anonymous):

28 +7/4r=-32 I need to see all the steps please

OpenStudy (anonymous):

First subtract 28 from both sides 28 + 7/4r = -32 -28 -28 7/4r = -50 Now multiply both sides by 4/7 4/7 * 7/4r = -50 * 4/7 r = -200/7

OpenStudy (anonymous):

I`m sorry,i think I wrote it out wrong.28 +7/4r , r=-32

OpenStudy (anonymous):